perform the indicated operations. write the resulting polynomial in standard form.
(-5x^6-13x^5-19) + (3x^6+9x^5+3)
a. -2x^6-4x^5-16
b. -2x^6+14x^5+22
c. -22x^11
d. -2x^6-4x^5+22

Answer by integral13(8)

-5x^6-13x^5-19
+ 3x^6+9x^5 +3
_________________________
Make sure your LIKE VARIABLES ARE ALL IN LINE!!!
-5x^6+3x^6=-2x^6
-13x^5+9x^5=-4x^5
-19+3= -16
Therefore, the answer is...
-2x^6-4x^5-16. (a.)
Hope this helps!!!
integral13